40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Apr 6 02:07:40 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source 
    Reply  Wed Apr 10 00:33:09 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source Modulation_frequency_100Hz.jpgModulation_frequency_30Hz.jpgModified_setup.JPG
       Reply  Wed Apr 10 16:58:54 2019, rana, Update, IOO, fiber MZ for NPRO freq noise measurements 
          Reply  Wed Apr 10 22:59:22 2019, gautam, Update, IOO, Spooled fiber 
             Reply  Thu Apr 11 09:05:06 2019, Anjali, Update, IOO, Spooled fiber Frequecy_modulation_32_Hz.pdfFrequecy_modulation_140_Hz.pdfFrequecy_modulation_300_Hz.pdfWithout_modulation.pdfNew_fiber_spool.JPG
       Reply  Fri Apr 12 01:22:27 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source 
          Reply  Mon Apr 15 22:39:10 2019, gautam, Update, Frequency noise measurement, Alternate setup with PSL pickoff 
             Reply  Wed Apr 17 00:43:38 2019, gautam, Update, Frequency noise measurement, MZ interferometer ---> DAQ 
                Reply  Thu Apr 25 03:32:25 2019, Anjali, Update, Frequency noise measurement, MZ interferometer ---> DAQ Time_domain_output.pdfFrequency_noise.pdfschematic_heterodyne_setup.pngNoise_budget_1_micron_in_Hz_per_rtHz.pdf
                   Reply  Thu Apr 25 10:25:19 2019, gautam, Update, Frequency noise measurement, Homodyne v Heterodyne 
                      Reply  Thu Apr 25 15:47:54 2019, Anjali, Update, Frequency noise measurement, Homodyne v Heterodyne Experimental_setup.JPG
Message ID: 14544     Entry time: Mon Apr 15 22:39:10 2019     In reply to: 14540     Reply to this: 14547
Author: gautam 
Type: Update 
Category: Frequency noise measurement 
Subject: Alternate setup with PSL pickoff 

[anjali, gautam]

just main points, anajli is going to fill out the details.

To rule out mode-matching as the reason for non-ideal output from the MZ, I suggested using the setup I have on the NW side of the PSL enclosure for the measurement. This uses two identical fiber collimators, and the distance between collimator and recombination BS is approximately the same, so the spatial modes should be pretty well matched. 

The spooled fiber we found was not suitable for use as it had a wide key connector and I couldn't find any wide-key FC/PC to narrow-key FC/APC adaptors. So we decided to give the fiber going to the Y end and back (~90m estimated length) a shot. We connected the two fibers at the EY table using a fiber mating sleeve (so the fiber usually bringing the IR pickoff from EY to the PSL table was disconnected from its collimator). 

In summary, we cannot explain why the contrast of the MZ is <5%. Spatial mode-overlap is definitely not to blame. Power asymmetry in the two arms of the MZ is one possible explanation, could also be unstable polarization, even though we think the entire fiber chain is PM. Anjali is investigating.

 


We saw today that the Thorlabs PM beam splitters (borrowed from Andrew until our AFW components arrive) do not treat the two special axes (fast and slow) of the fiber on equal footing. When we coupled light into the fast axis, we saw huge asymmetry between the two split arms of the beamsplitter (3:1 ratio in power instead of the expected 1:1 for a 50/50 BS). Looking at the patch cord with an IR viewer, we also saw light leaking through the core along it. Turns out this part is meant to be used with light coupled to the slow axis only.

ELOG V3.1.3-